# Knowledge Graphs for the Life Sciences: Recent Developments, Challenges and Opportunities

### v1

@@ 4. Life Science Knowledge Discovery
Knowledge Graphs are powerful tools for representing complex biomedical knowledge, including molecular interactions, signalling pathways, disease co-morbidities, and more. Overviews of graph representation learning in biomedicine for healthcare applications and polypharmacy tasks are presented in [110] and [54] respectively. In graph representation learning, the graph’s topology is leveraged to create compact vector embeddings. Through nonlinear transformations, high-dimensional information about a node’s graph neighborhood is distilled into low-dimensional vectors, where similar nodes are embedded close together in the vectorial space. Embeddings have been shown to be valuable for handling numerous relations in a KG while efficiently exploiting relation sparsity using vector computations. These optimized representations are subsequently used to train downstream models for various tasks, such as predicting property values of specific nodes (e.g., protein function), predicting links between nodes (e.g., binding affinity between molecules and protein targets), or performing classification tasks (e.g., predicting the toxicity profile of a candidate drug, or risk of readmission for a patient).
It is worth mentioning that among the existing works for life science knowledge discovery, different kinds of KGs have been exploited. The schema-less KG can be used to model different kinds of interaction between instances such as proteins and drugs; the taxonomy alike simple ontology is often used to represent concepts and their hierarchy such as protein functions defined in the gene ontology, chemical compounds, species, and diseases; expressive OWL ontologies and schema-based KGs can be used to model complex logical relationships between concepts, besides simple interaction between instances. Such diverse knowledge representation capabilities make KGs more flexible in modeling the input data and prediction targets of different knowledge discovery tasks, than graphs and tabular data that are widely used in previous pure machine learning-based methods.
In the following, we present some typical use cases, where machine learning techniques (including graph representation learning and language models) are applied over KGs built from diverse sources and domain ontologies, to facilitate life science discovery.
4.1. What has been done: use cases and their recent developments
Therapeutics and Drug Discovery: Learning a representation using multi-modal and heterogeneous knowledge. Drug discovery entails exploring an extremely large space of potential drug candidates. AI can help to accelerate this process by narrowing down the most promising candidates before expensive experimentation. The key to leveraging predictive and generative models for candidate solution generation lies in learning an effective multi-modal representation of protein targets, molecules and diseases among others. Recent research has focused on applying language models over large databases of proteins or molecules for self-supervised representation learning, such as ESM [151] and ProteinBERT [11] for protein sequences, or Molformer for the molecule simplified molecular-input line-entry system (SMILES) [154]. These models have exhibited remarkable success in tasks such as predicting protein interactions, binding affinity between drugs and targets, and protein functions and structures. However, these existing pre-trained sequence-based models often neglect to incorporate background knowledge from diverse sources, for example, biological structural knowledge.
Nonetheless, recent research indicates that incorporating existing expressive factual knowledge can improve results in downstream machine learning tasks. To enhance Protein Language Models (PLM), approaches such as OntoProtein [195] and KeAP [197] use a KG of protein sequences augmented with textual annotations from the Gene Ontology (GO). OntoProtein was the first to inject gene ontology descriptions into a PLM for sequences to predict protein interactions, function and contact prediction. OntoProtein proposes to reconstruct masked amino acids while minimizing the embedding distance between the contextual representation of proteins and associated knowledge terms. Similarly, ProtST [189] uses a dataset of protein sequences augmented with textual property descriptions from biomedical texts and jointly trains a PLM with a biomedical language model.
Knowledge Graphs are suitable data models for expressing heterogeneous knowledge and facilitating end-to-end learning [184]. An entity in a KG can have multiple attributes with different modalities - where each modality provides extra information about the entity - as well as relations to and from entities in other sources. Graph Neural Networks (GNN) have been used to capture inter-dependencies and diverse types of interactions between heterogeneous entity types and multimodal attributes in KGs [103]. They achieve this by iteratively aggregating information from neighbouring nodes (through a process called message passing) and employing scoring functions to optimize the learned embeddings for downstream tasks. Otter-Knowledge [103] incorporates a heterogeneous KG (schema-based, containing concepts and their attributes) from diverse sources and modalities, i.e., each node has a particular mode that qualifies its type (text, image, protein sequence, molecule, etc.) and initial embeddings for each node are computed based on their modality. A GNN is then used to enrich protein and molecule representations and train a model to produce final node embeddings. The model is able to produce representations for entities that were not seen during training and achieve state-of-the-art results in the Therapeutic Data Commons (TDC) benchmarks [75] for drug-target binding affinity prediction. TxGNN [74] uses a GNN pre-trained on a large heterogeneous, multi-relational KG of diseases and therapeutic candidates constructed from various knowledge bases. TxGNN obtains a signature vector for each disease based on its neighboring proteins, exposure and other biomedical entities to compute a disease similarity and predict drug indication/contraindication for poorly characterized diseases.
Protein Function Prediction with the Gene Ontology. Conducting physical experiments for identifying protein functions is time and resource consuming. With the development of machine learning, protein function prediction (which is the task of predicting a given protein with multiple and potentially hierarchical classes – functions – defined in GO) has been widely investigated in recent years [196, 174]. A large part of these works such as GOLabler [192] focus on exploring feature extraction, feature ensemble, and automatic feature learning of the proteins. For example, GOLabler [192] utilizes five kinds of different protein sequence information while DeepGraphGO [191] builds a network of proteins and learns protein features via a Graph Neural Network. Recent methods attempt to further exploit inter-function (class) relationships that are defined in GO for better performance. For example, DeepGOZero [102] and HMI [188] use formal semantics including the class hierarchy, class disjointness axioms and complex class restrictions in OWL as additional constraints for training the multi-label classifier for protein function prediction. Protein function prediction is a representative multi-label classification problem where complex relationships of the labels are defined in a KG and can be used for performance augmentation. It is quite common in machine learning applications in the life sciences, such as the above mentioned automated clinical coding where the codes’ semantics are modeled by the ICD ontology, and ecotoxicological effect prediction where the multiple affected species of a chemical to predict form a taxonomy.
Predictions for Healthcare using Ontologies with Clinical Data. Digital Healthcare involves predictions using clinical data and ontologies, including diagnosis (e.g., rare diseases) and procedure predictions (e.g., ICU readmissions). A related concept is personalized medicine, which is achieved through the matching and fusion of knowledge from diverse sources, and plays a significant role in the prediction tasks. This often involves matching multiple ontologies [159], integrating curated databases (e.g., pharmacogenomics, molecules and proteins knowledge bases), mining knowledge from scientific literature [187] and person-centered clinical knowledge extracted from EHR or claim data, with distinguishing risk factors or cohorts’ demographics (e.g., age and gender), which could enhance predictions related to adverse effects [126] or rare diseases for which there are not enough labeled datasets [2]. For example, SHEPHERD [2] incorporates a multi-relational KG (extracted from PrimeKG [20]) of diseases, phenotypes and genes, and leverages patient simulated data to discover novel connections between patients’ clinical, phenotype and gene information to accelerate the diagnoses of rare diseases. Knowledge-guided learning is achieved by training a GNN to represent each patient’s subgraphs of phenotypes in relation to other gene, phenotype, and disease associations within the KG, such that embeddings are informed by all of the existing biomedical knowledge captured in the network topology.
The approach in [16] constructs a KG (using expressive OWL ontologies) to predict ICU (intensive care units) readmission risk by enriching EHR data with semantic annotations from various biomedical ontologies in BioPortal. These predictions are based on KG embedding, such as RDF2vec, OPA2vec, and TransE, and classical machine learning methods, such as Logistic Regression, Random Forest, Naive Bayes and Support Vector Machines. Drawing from the Health & Social Person-centric Ontology (HSPO) [168], which focuses on multiple clinical, social and demographic facets for a patient or cohort, the approach presented in [169] builds a person-centric KG (expressive OWL ontology with TBox and ABox) from structured and unstructured data in EHR). Subsequently, a representation learning approach using GNNs is used to predict readmissions to the ICU.
4.2. What are the challenges?
@@
  • How to create an efficient multi-modal representation of knowledge to enable discovery? Most current state-of-the-art methods build learned graph representations based on isolated modalities. Multimodal KGs can explicitly capture labelled nodes and edges, each with well-defined meanings, across heterogeneous node types, relations and modalities (such as text, images, protein sequences, molecules fingerprints, diseases and more) [20, 103]. Incorporating KGs with multiple modalities for representation learning requires computationally scalable methods to compute the initial embeddings for each modality, as a preliminary step to learn computable representations of large knowledge. Furthermore, robust learning techniques are needed for generalizing the learned representations to nodes with unseen or missing modalities, thereby enabling the discovery of new knowledge. An example would be inferring properties of proteins for which only the sequence is known.
@@
  • How to efficiently utilize and fuse heterogeneous datasets, such as human-curated domain knowledge bases, scientific literature and person-centered health records, for knowledge discovery? State of the art shows that representations can be enhanced by incorporating richer information available across different sources [74, 103, 159]. Bringing in more data during training is needed to learn representations that can be applied to a broader range of downstream prediction tasks. However, learning from large and diverse KGs requires addressing challenges such as alignment, noise handling, balancing rich expressive knowledge with scalability and dealing with knowledge inconsistency. Moreover, more robust learning methods are needed for generalizing the learned representation to multiple downstream tasks (e.g., knowledge-aware transfer, zero-shot and few-shot learning [23]). An important aspect in this regard is addressing the disparity between all of the knowledge accessible during pre-training and the knowledge accessible or relevant for the downstream fine-tuning [74, 103].
5. Knowledge Graphs for Explainable AI
Machine Learning (ML) and Artificial Intelligence (AI) methods are widely employed to tackle complex problems in many domains, including life sciences such as chemistry or biomedicine. Yet many of those methods operate as a “black-box”, not enabling domain experts to understand the reasoning behind their predictions [93]. This is a major concern, especially for applications in areas with a potential impact on human lives, or areas with legally enforced accountability or transparency [146]. Moreover, understanding the workings of AI methods is also crucial in the context of scientific applications, such as those described in Section 4, where explaining the prediction process can help elucidate natural phenomena [42].
One way to address this issue is to employ the methods of eXplainable Artificial Intelligence (XAI). Although this is a topic long explored in the AI research community, there is still no widely-accepted definition of explainability, with many terms being used interchangeably, such as interpretability, comprehensibility, understandability and transparency [8]. Barredo et al. define explainability as the ability of a model to make its functioning clearer to an audience [8]. A slightly different definition is given in the previous survey [57]: “an interface between humans and a decision maker that is at the same time both an accurate proxy of the decision maker and comprehensible to humans”. Both definitions focus on the audience, for whom is the model explainable, but the second suggests an explanation is another artefact produced by a model or alongside the model.
There are two distinguishable audiences in the context of the life sciences: scientists (researchers) and healthcare practitioners [170]. For the first group, the explanation is used as a guide to understanding within life sciences research for scientific discovery. As a result, the explanation may exist in a well-bounded context of a hypothesis or research project. On the other hand, practitioners are involved directly in decisions with impact on healthcare. They need to consider the output of the model in an open context, and sometimes also to explain the output to a patient who is not a domain expert.
A number of approaches for XAI emerge from the literature and broadly contain two parts: (1) transparent box design, which includes algorithms such as decision trees, where models can be directly interpreted by users and therefore an explanation of an output results in simply following the decision paths that relate input to output; (2) post hoc interpretability, which provides an explanation to a black-box model using additional methods such as probing, perturbing, or by constructing surrogate models for general ML or AI methods [93, 170].
Utilization of KGs can greatly enhance XAI qualities as KGs are ideal for improving the model’s interpretability, explainability, and understandability. Some methods are directly built around KGs and thus take full advantage of them. Examples of those methods may include methods that are using paths [164], predicting links, or performing reasoning [34]. Other methods can be enhanced using the KG (e.g., [129]). Yet the enhancement effect greatly depends on the place where KGs are employed and iteratively applied: pre-model (e.g., KG construction, potentially multi-modal), in-model (e.g., integrating KG with machine learning models), and post-model (e.g., reviewing and updating KG by domain experts to be applied in the next iteration to enhance machine learning models and their explanability) [146]. For example in in-model use, a model can be pre-trained using a KG, and an example of a pre-trained language model is SapBERT [111], which utilises synonyms in the UMLS Metathesaurus to further pre-train a BERT language model. This can not only be beneficial for performance [195], but can also potentially enhance post-model explanation since the trained features are aligned with the KG [146].
5.1. What has been done: use cases and recent developments
Explainable AI for Healthcare Practice. The utilization of AI in healthcare practice raises the concern of leaving life-critical decisions to black-box models [146, 170]. For example, in the field of precision medicine which aims at tailoring drug treatments and dosages to each patient, clinicians require more information from a model than a simple binary decision [8]. The interpretability and explainability of AI models is thus an essential characteristic to make outputs understandable and transparent. This would enforce both clinicians’ and patients’ trust in models by complementing (and not substituting) clinicians’ explanations [21, 146, 170].
To illustrate, this direction has been envisioned for several healthcare scenarios. Explainable AI models could support the experts in finding clinical trials that are appropriate based on patient history [170]. Counterintuitive or unreliable predictions that could have serious consequences could be explained, and thus prevented [170, 15, 92]. Some also envision such models to be used to explain and debunk healthcare-related misinformation [146]. As aforementioned, it is noteworthy that different kinds of explanations should be employed depending on the target audience, e.g., scientific explanations for evidence or trace-based explanations for treatment [21].

### v4

@@ 4 Life Science Knowledge Discovery
Research into AI technologies – including machine learning and KG-based reasoning – to accelerate the pace of scientific discovery is an emerging and rapidly developing field. The challenge lies in assisting scientists to uncover new knowledge and solutions, such as discovering novel therapeutic opportunities, identifying candidate molecular drugs to treat complex diseases or alternatively new uses for existing drugs, and supporting more personalized predictions. Knowledge Graphs are powerful tools for representing complex biomedical knowledge, including molecular interactions, signaling pathways, disease co-morbidities, and more. Overviews of graph representation learning in biomedicine for healthcare applications and polypharmacy tasks are presented in [107] and [52] respectively. In graph representation learning, the graph’s topology is leveraged to create compact vector embeddings. Through nonlinear transformations, high-dimensional information about a node’s graph neighborhood is distilled into low-dimensional vectors, where similar nodes are embedded close together in the vectorial space. Embeddings have been shown to be valuable for handling numerous relations in a KG while efficiently exploiting relation sparsity using vector computations. These optimized representations are subsequently used to train downstream models for various tasks, such as predicting property values of specific nodes (e.g., protein function), predicting links between nodes (e.g., binding affinity between molecules and protein targets), or performing classification tasks (e.g., predicting the toxicity profile of a candidate drug, or risk of readmission for a patient). It is worth mentioning that among the existing works for life science knowledge discovery, different kinds of KGs have been exploited. The schema-less KG can be used to model different kinds of interaction between instances such as proteins and drugs; the taxonomy alike simple ontology is often used to represent concepts and their hierarchy such as protein functions defined in the gene ontology, chemical compounds, species, and diseases; expressive OWL ontologies and schema-based KGs can be used to model complex logical relationships between concepts, besides simple interaction between instances. Such diverse knowledge representation capabilities make KGs more flexible in modeling the input data and prediction targets of different knowledge discovery tasks, than graphs and tabular data that are widely used in previous pure machine learning-based methods. In the following, we present some typical use cases, where machine learning techniques (including graph representation learning and language models) are applied over KGs built from diverse sources and domain ontologies, to facilitate life science discovery.
4.1 What has been done: use cases and their recent developments
Therapeutics and Drug Discovery: Learning a representation using multi-modal and heterogeneous knowledge. Drug discovery entails exploring an extremely large space of potential drug candidates. AI can help to accelerate this process by narrowing down the most promising candidates before expensive experimentation. The key to leveraging predictive and generative models for candidate solution generation lies in learning an effective multi-modal representation of protein targets, molecules and diseases among others. Recent research has focused on applying language models over large databases of proteins or molecules for self-supervised representation learning, such as ESM [148] and ProteinBERT [10] for protein sequences, or Molformer for the molecule simplified molecular-input line-entry system (SMILES) [151]. These models have exhibited remarkable success in tasks such as predicting protein interactions, binding affinity between drugs and targets, and protein functions and structures. However, these existing pre-trained sequence-based models often neglect to incorporate background knowledge from diverse sources, for example, biological structural knowledge. Nonetheless, recent research indicates that incorporating existing expressive factual knowledge can improve results in downstream machine learning tasks. To enhance Protein Language Models (PLM), approaches such as OntoProtein [192] and KeAP [194] use a KG of protein sequences augmented with textual annotations from the Gene Ontology (GO). OntoProtein was the first to inject gene ontology descriptions into a PLM for sequences to predict protein interactions, function and contact prediction. OntoProtein proposes to reconstruct masked amino acids while minimizing the embedding distance between the contextual representation of proteins and associated knowledge terms. Similarly, ProtST [186] uses a dataset of protein sequences augmented with textual property descriptions from biomedical texts and jointly trains a PLM with a biomedical language model. Knowledge Graphs are suitable data models for expressing heterogeneous knowledge and facilitating end-to-end learning [181]. An entity in a KG can have multiple attributes with different modalities - where each modality provides extra information about the entity - as well as relations to and from entities in other sources. Graph Neural Networks (GNN) have been used to capture inter-dependencies and diverse types of interactions between heterogeneous entity types and multimodal attributes in KGs [100]. They achieve this by iteratively aggregating information from neighboring nodes (through a process called message passing) and employing scoring functions to optimize the learned embeddings for downstream tasks. Otter-Knowledge [100] incorporates a heterogeneous KG (schema-based, containing concepts and their attributes) from diverse sources and modalities, i.e., each node has a particular mode that qualifies its type (text, image, protein sequence, molecule, etc.) and initial embeddings for each node are computed based on their modality. A GNN is then used to enrich protein and molecule representations and train a model to produce final node embeddings. The model is able to produce representations for entities that were not seen during training and achieve state-of-the-art results in the Therapeutic Data Commons (TDC) benchmarks [73] for drug-target binding affinity prediction. TxGNN [72] uses a GNN pre-trained on a large heterogeneous, multi-relational KG of diseases and therapeutic candidates constructed from various knowledge bases. TxGNN obtains a signature vector for each disease based on its neighboring proteins, exposure and other biomedical entities to compute a disease similarity and predict drug indication/contraindication for poorly characterized diseases. Protein Function Prediction with the Gene Ontology. Conducting physical experiments for identifying protein functions is time and resource consuming. With the development of machine learning, protein function prediction (which is the task of predicting a given protein with multiple and potentially hierarchical classes – functions – defined in GO) has been widely investigated in recent years [193, 171]. A large part of these works such as GOLabler [189] focus on exploring feature extraction, feature ensemble, and automatic feature learning of the proteins. For example, GOLabler [189] utilizes five kinds of different protein sequence information while DeepGraphGO [188] builds a network of proteins and learns protein features via a Graph Neural Network. Recent methods attempt to further exploit inter-function (class) relationships that are defined in GO for better performance. For example, DeepGOZero [99] and HMI [185] use formal semantics including the class hierarchy, class disjointness axioms and complex class restrictions in OWL as additional constraints for training the multi-label classifier for protein function prediction. Protein function prediction is a representative multi-label classification problem where complex relationships of the labels are defined in a KG and can be used for performance augmentation. It is quite common in machine learning applications in the life sciences, such as the above mentioned automated clinical coding where the codes’ semantics are modeled by the ICD ontology, and ecotoxicological effect prediction where the multiple affected species of a chemical to predict form a taxonomy. Predictions for Healthcare using Ontologies with Clinical Data. Digital Healthcare involves predictions using clinical data and ontologies, including diagnosis (e.g., rare diseases) and procedure predictions (e.g., ICU readmissions). A related concept is personalized medicine, which is achieved through the matching and fusion of knowledge from diverse sources, and plays a significant role in the prediction tasks. This often involves matching multiple ontologies [156], integrating curated databases (e.g., pharmacogenomics, molecules and proteins knowledge bases), mining knowledge from scientific literature [184] and person-centered clinical knowledge extracted from EHR or claim data, with distinguishing risk factors or cohorts’ demographics (e.g., age and gender), which could enhance predictions related to adverse effects [123] or rare diseases for which there are not enough labeled datasets [2]. For example, SHEPHERD [2] incorporates a multi-relational KG (extracted from PrimeKG [19]) of diseases, phenotypes and genes, and leverages patient simulated data to discover novel connections between patients’ clinical, phenotype and gene information to accelerate the diagnoses of rare diseases. Knowledge-guided learning is achieved by training a GNN to represent each patient’s subgraphs of phenotypes in relation to other gene, phenotype, and disease associations within the KG, such that embeddings are informed by all of the existing biomedical knowledge captured in the network topology. The approach in [15] constructs a KG (using expressive OWL ontologies) to predict ICU (intensive care units) readmission risk by enriching EHR data with semantic annotations from various biomedical ontologies in BioPortal. These predictions are based on KG embedding, such as RDF2vec, OPA2vec, and TransE, and classical machine learning methods, such as Logistic Regression, Random Forest, Naive Bayes and Support Vector Machines. Drawing from the Health & Social Person-centric Ontology (HSPO) [165], which focuses on multiple clinical, social and demographic facets for a patient or cohort, the approach presented in [166] builds a person-centric KG (expressive OWL ontology with TBox and ABox) from structured and unstructured data in EHR). Subsequently, a representation learning approach using GNNs is used to predict readmissions to the ICU.
4.2 What are the challenges?
@@
  • How to create an efficient multi-modal representation of knowledge to enable discovery? Most current state-of-the-art methods build learned graph representations based on isolated modalities. Multimodal KGs can explicitly capture labeled nodes and edges, each with well-defined meanings, across heterogeneous node types, relations and modalities (such as text, images, protein sequences, molecules fingerprints, diseases and more) [19, 100]. Incorporating KGs with multiple modalities for representation learning requires computationally scalable methods to compute the initial embeddings for each modality, as a preliminary step to learn computable representations of large knowledge. Furthermore, robust learning techniques are needed for generalizing the learned representations to nodes with unseen or missing modalities, thereby enabling the discovery of new knowledge. An example would be inferring properties of proteins for which only the sequence is known.
@@
  • How to efficiently utilize and fuse heterogeneous datasets, such as human-curated domain knowledge bases, scientific literature and person-centered health records, for knowledge discovery? State of the art shows that representations can be enhanced by incorporating richer information available across different sources [72, 100, 156]. Bringing in more data during training is needed to learn representations that can be applied to a broader range of downstream prediction tasks. However, learning from large and diverse KGs requires addressing challenges such as alignment, noise handling, balancing rich expressive knowledge with scalability and dealing with knowledge inconsistency. Moreover, more robust learning methods are needed for generalizing the learned representation to multiple downstream tasks (e.g., knowledge-aware transfer, zero-shot and few-shot learning [22]). An important aspect in this regard is addressing the disparity between all of the knowledge accessible during pre-training and the knowledge accessible or relevant for downstream fine-tuning [72, 100].
5 Knowledge Graphs for Explainable AI
Machine Learning (ML) and Artificial Intelligence (AI) methods are widely employed to tackle complex problems in many domains, including life sciences such as chemistry or biomedicine. Yet many of those methods operate as a “black-box”, not enabling domain experts to understand the reasoning behind their predictions [90]. This is a major concern, especially for applications in areas with a potential impact on human lives, or areas with legally enforced accountability or transparency [143]. Moreover, understanding the workings of AI methods is also crucial in the context of scientific applications, such as those described in Section 4, where explaining the prediction process can help elucidate natural phenomena [40]. One way to address this issue is to employ the methods of eXplainable Artificial Intelligence (XAI). Although this is a topic long explored in the AI research community, there is still no widely-accepted definition of explainability, with many terms being used interchangeably, such as interpretability, comprehensibility, understandability and transparency [7]. Barredo et al. define explainability as the ability of a model to make its functioning clearer to an audience [7]. A slightly different definition is given in the previous survey [55]: “an interface between humans and a decision maker that is at the same time both an accurate proxy of the decision maker and comprehensible to humans”. Both definitions focus on the audience, for whom is the model explainable, but the second suggests an explanation is another artifact produced by a model or alongside the model. There are two distinguishable audiences in the context of the life sciences: scientists (researchers) and healthcare practitioners [167]. For the first group, the explanation is used as a guide to understanding within life sciences research for scientific discovery. As a result, the explanation may exist in a well-bounded context of a hypothesis or research project. On the other hand, practitioners are involved directly in decisions with impact on healthcare. They need to consider the output of the model in an open context, and sometimes also to explain the output to a patient who is not a domain expert. A number of approaches for XAI emerge from the literature and broadly contain two parts: (1) transparent box design, which includes algorithms such as decision trees, where models can be directly interpreted by users and therefore an explanation of an output results in simply following the decision paths that relate input to output; (2) post hoc interpretability, which provides an explanation to a black-box model using additional methods such as probing, perturbing, or by constructing surrogate models for general ML or AI methods [90, 167]. Utilization of KGs can greatly enhance XAI qualities as KGs are ideal for improving the model’s interpretability, explainability, and understandability. Some methods are directly built around KGs and thus take full advantage of them. Examples of those methods may include methods that are using paths [161], predicting links, or performing reasoning [33]. Other methods can be enhanced using the KG (e.g., [126]). Yet the enhancement effect greatly depends on the place where KGs are employed and iteratively applied: pre-model (e.g., KG construction, potentially multi-modal), in-model (e.g., integrating KG with machine learning models), and post-model (e.g., reviewing and updating KG by domain experts to be applied in the next iteration to enhance machine learning models and their explanability) [143]. For example in in-model use, a model can be pre-trained using a KG, and an example of a pre-trained language model is SapBERT [108], which utilizes synonyms in the UMLS Metathesaurus to further pre-train a BERT language model. This can not only be beneficial for performance [192], but can also potentially enhance post-model explanation since the trained features are aligned with the KG [143].
5.1 What has been done: use cases and recent developments